OpenStudy (anonymous):

What is the function that describes this sequence? x y 1 3 2 6 3 9 4 12 A. Y=3/N B. Y=N/3 C. Y=3n D. Y=n + 3

OpenStudy (experimentx):

all multiplied by 3

OpenStudy (anonymous):

C. Y=3N ^_^ 1*3=3 2*3=6 3*3=9 4*3=12

OpenStudy (anonymous):

Thanks! was having a little trouble with this one, and did not catch onto it until @experimentX said something about it!